# Transverse vortices induced by modulated granular shear flows of elongated particles

### v1

@@ VII Conclusion and perspectives
We have studied the response of a granular shear flow of elongated particles to a perturbative streamwise forcing, running DEM numerical simulations and analytically computing the linear flow response from the granular rheology generalized to this type of grain. Both reveal the emergence of transverse secondary flows that, upon averaging over sufficiently long times, generate a four-vortex pattern when the perturbation is a single Fourier mode. A central result is the observation and understanding of the inversion of this pattern when the cell aspect ratio is varied. We also observe a pattern inversion when the particle contact friction is changed at a fixed cell aspect ratio but for a limited range.
@@
We have discussed the possible reasons why the agreement between simulations and theory is qualitative only. This calls for longer simulations with a smaller perturbation amplitude or for runs with larger systems in order to check finite-size effects. In particular, we expect that gradients that are too strong at the scale of a few grains would not match well with the continuum theoretical approach and/or could be associated with non-local effects, which the present theory does not account for. Such tests are however computationally expensive, as a long time average is necessary to see the emergence of these vortex patterns out of the noise. We leave more systematic studies with such additional simulations for later work.
@@
Finally, this work offers a new rheological way to look at the heaping problem Wortel et al. (2015); Fischer et al. (2016). This phenomenon, which also shows secondary vortex flows, is of course more complicated than what we observe here in the perturbed shear cell configuration. Not only the split-bottom geometry is less easy to handle, but also because gravity is present and can affect these flows through buoyancy convection associated with spatial variations of the packing fraction. Nevertheless, it is striking that the amplitude of this heaping is sensitive to the filling height, i.e. the equivalent of the cell aspect ratio here. Also, the understanding that varying a microscopic parameter, such as the contact friction coefficient $\mu_{p}$, can inverse the vortex pattern under some conditions provides a possible interpretation of the observation that non-convex granular particles generate dipping rather than heaping in the split-bottom cell Mohammadi et al. (2022). Clearly, more numerical simulations as well as experiments with different grains and geometries to investigate the effect of varied normal stress differences would be very instructive.
@@ Acknowledgments
The authors thank Dániel Bálint Nagy for the input related to the generalized granular rheology. We also thank Pierre Jop, Cyril Gadal, and Julian Talbot for useful discussions. We acknowledge support from CNRS PICS Grant No. 08187, the Hungarian Academy of Sciences Grant No. NKM-2018-5, the COST action ”ON-DEM” (No. CA22132), and the Spanish Government through grant No. PID2023-146422NB-I00 supported by MICIU/AEI/10.13039/501100011033.
@@ Appendix A Rheological parameters
Nagy et al. Nagy et al. (2020) studied the rheological parameters, like the effective friction coefficient ($\mu$) and the normal stress differences ($N_{1}$ and $N_{2}$) of elongated grains in a plane Couette flow for various particle aspect ratios ($\alpha$), particle friction coefficients ($\mu_{p}$), and inertial numbers ($I$). In Fig. A1 we present their measurements for the case of $I=10^{-2}$, and the interpolated values of the rheological parameters as functions of particle friction coefficient (panels a-c) and aspect ratio (panels d-f). We remind the reader that $\nu_{1}=\frac{N_{1}+N_{2}}{\Pi}$, $\nu_{2}=\frac{N_{2}}{\Pi}$, and $\pi=1-\tfrac{1}{3}(\nu_{1}+\nu_{2})$. Using these results we calculated the linear response of our system ($a_{y\text{,rescaled}}$) which is presented in Fig. A2.
Figure A1: $\mu$ (a,d), $\nu_{2}$ (b,e), and $\pi$ (c,f) as a function of the particle-particle coefficient of friction ($\mu_{p}$) (a-c) and the particle aspect ratio ($\alpha$) (d-f) for cell aspect ratio $L_{z}/L_{y}{=}1$ and inertial number $I=10^{-2}$ using results of Nagy et al. Nagy et al. (2020).
Figure A2: The rescaled response of the system ($a_{y\text{,rescaled}}{=}a_{y}(L_{y}^{2}{+}L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$) as function of the particle-particle coefficient of friction ($\mu_{p}$) (a) and the particle aspect ratio ($\alpha$) (b) for cell aspect ratio $L_{z}/L_{y}{=}1$ (continuous lines), and $I=10^{-2}$. In panel (a) for $\alpha{=}1.5$ we show the influence of a slight change (±$\pm 10$%) of the cell aspect ratio on the response of the system (dashed green lines). For cell aspect ratios slightly larger than $1$ we have a change of sign in the response.
@@ Appendix B Time evolution of the secondary flow pattern
To show the time evolution of the secondary flow vortex pattern, in the Supplemental Material (SM) we present the time evolution of the velocity fields for two cell aspect ratios presented in Fig. 3, for $L_{z}/L_{y}\approx 0.5$ (SM Video 1, SM0 (a)) and $L_{z}/L_{y}\approx 1$ (SM Video 2, SM0 (b)). The flow pattern with four vortices starts to form relatively quickly; however, the strength and relative position of the vortices change a lot as the system evolves.

### v2

@@ VII Conclusion and perspectives
We studied the response of a granular shear flow of elongated particles to a perturbative streamwise forcing, running DEM numerical simulations and analytically computing the linear flow response from the granular rheology generalized to this type of grain. Both reveal the emergence of transverse secondary flows that, upon averaging over sufficiently long times, generate a four-vortex pattern when the perturbation is a single Fourier mode. A central result is the observation and understanding of the inversion of this pattern when the cell aspect ratio is varied. We also observe a pattern inversion when the particle contact friction is changed at a fixed cell aspect ratio, but for a limited range.
@@
We discussed the possible reasons why the agreement between simulations and theory is qualitative only. This calls for longer simulations with a smaller perturbation amplitude or for runs with larger systems to check finite-size effects. In particular, we expect that gradients that are too strong at the scale of a few grains would not match well with the continuum theoretical approach and/or could be associated with nonlocal effects, which the present theory does not account for. Such tests are, however, computationally expensive, as a long time average is necessary to see the emergence of these vortex patterns out of the noise. We leave more systematic studies with such additional simulations for later work.
@@
Finally, this work offers additional insight into the heaping problem observed in split-bottom shear devices Wortel et al. (2015); Fischer et al. (2016). This phenomenon, which also shows secondary vortex flows, is, of course, more complicated than what we observe here in the perturbed shear cell configuration. Not only is the split-bottom geometry less easy to handle, but also because gravity is present and can affect these flows through buoyancy convection associated with spatial variations of the packing fraction. Nevertheless, it is striking that the amplitude of this heaping is sensitive to the filling height, i.e., the equivalent of the cell aspect ratio here. Also, the understanding that varying a microscopic parameter, such as the contact friction coefficient $\mu_{p}$, can invert the vortex pattern under some conditions provides a possible interpretation of the observation that nonconvex granular particles generate dipping rather than heaping in the split-bottom cell Mohammadi et al. (2022). Clearly, more numerical simulations as well as experiments with different grains and geometries to investigate the effect of varied normal stress differences would be very instructive.
Figure 7: Response of the system in the shear direction for the same simulations as presented in Fig. 4. The response is rescaled by the system size ($a_{x\text{,rescaled}}=a_{x}/L_{z}^{2}$), and plotted as a function of cell aspect ratio for two particle aspect ratios (a) $\alpha=1.5$ and (b) $\alpha=2$, and for different particle coefficients of friction $\mu_{p}$ (legends). The strength of perturbation is $A=2\times 10^{-4}$. Dashed lines show the expected values according to the theory. The symbols show the results of simulations, averaged over five different realizations. Their span is in the order of the point size. Importantly, note the different vertical axes on these graphs for simulation results (left) and theoretical predictions (right) — a qualitative-only agreement is reached.
@@ Acknowledgments
The authors thank Dániel Bálint Nagy for the input related to the generalized granular rheology. We also thank Pierre Jop, Cyril Gadal, and Julian Talbot for useful discussions. We acknowledge support from CNRS PICS/IEA Grant No. 08187, the Hungarian Academy of Sciences Grant No. NKM-2018-5, the COST action ”ON-DEM” (Grant No. CA22132), and the Spanish Government through Grant No. PID2023-146422NB-I00 supported by MICIU/AEI/10.13039/501100011033 and FEDER, EU.
Data availability
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
@@ Appendix A Rheological parameters
Nagy et al. Nagy et al. (2020) studied the rheological parameters, like the effective friction coefficient ($\mu$) and the normal stress differences ($N_{1}$ and $N_{2}$) of elongated grains in a plane Couette flow for various particle aspect ratios ($\alpha$), particle friction coefficients ($\mu_{p}$), and inertial numbers ($I$). In Fig. 5 we present their measurements for the case of $I=10^{-2}$, and the interpolated values of the rheological parameters as functions of particle friction coefficient [Figs. 5(a) to 5(c)] and aspect ratio [Figs. 5(d) to 5(f)]. We remind the reader that $\nu_{1}=\frac{N_{1}+N_{2}}{\Pi}$, $\nu_{2}=\frac{N_{2}}{\Pi}$, and $\pi=1-\tfrac{1}{3}(\nu_{1}+\nu_{2})$. Using these results, we calculated the linear response of our system ($a_{y\text{,rescaled}}$), which is presented in Fig. 6.
@@ Appendix B Time evolution of the secondary flow pattern
To show the time evolution of the secondary flow vortex pattern, in the Supplemental Material (SM) SM we present the time evolution of the velocity fields for two cell aspect ratios presented in Fig. 3(a), for $L_{z}/L_{y}\approx 0.5$ (SM Video 1, SM ) and $L_{z}/L_{y}\approx 1$ (SM Video 2, SM ). The flow pattern with four vortices starts to form relatively quickly; however, the strength and relative position of the vortices change a lot as the system evolves.